A kinetic sculpture team took
5
5 minutes to race down a river with a current of
4
4 mph, and
13
13 minutes to travel back. Use the following system of equations to find the rate
r
r of the kinetic sculpture team.
{
5

(
r
+
4
)
=
d
13

(
r

4
)
=
d
{5⋅(r+4)=d13⋅(r-4)=d

Answer by jorel555(1290)

4 mph=1/15 mile per minute
5(d+1/15)=13(d-1/15)
5d+5/15=13d-13/15
Multiply through by 15, we get:
75d+5=195d-13
120d=18
d=18/120=.15 miles per minute, or 9 miles per hour in still water!!!!!!!